# MagicDrive3D: Controllable 3D Generation for Any-View Rendering in Street Scenes

### v2

@@ 4.3 Enhanced Gaussian Splatting for Generated Content
Consequently, $\boldsymbol{\mu}^{o}_{p}(t)$ can manage the local dynamics driven by pixel-level disparities, while $\boldsymbol{\mu}$ focuses on the global geometric correlations. It ensures the quality of scene reconstruction by leveraging consistent parts across different viewpoints, simultaneously eliminating artifacts. Besides, with the analytical gradient w.r.t. $SE(3)$ pose of cameras (Matsuki et al., 2024), we also make the camera pose optimizable in the final few steps of GS iterations, which helps to mitigate the local dynamic from camera poses.
@@
Loss: Aligning Exposure with Appearance Modeling. Typical street view dataset is collected with multiple cameras, which capture views independently through auto-exposure and auto-white-balance (Caesar et al., 2020). Since the video generation is optimized to match the original data distribution, the differences from different cameras also exist in the generated data. The appearance differences are well-known issues for in-the-wild reconstruction (Martin-Brualla et al., 2021). In this paper, we propose a dedicated appearance modeling technique for GS representation.
We hypothesize that the disparity between different views can be represented by affine transformations $\mathbf{A}_{i}(\cdot)$ for $i$-th camera view. An Appearance Embedding (AE) map $\mathbf{e}_{i}\in\mathbb{R}^{w_{e}\times h_{e}\times c_{e}}$ is allocated for each view, and a Convolutional Neural Network (CNN) is utilized to approximate this transformation matrix $w_{\mathbf{A}}\in\mathbb{R}^{w\times h\times 3}$ (Appendix D contains more details). The final computation of the pixel-wise $\ell_{1}$ loss is conducted using the transformed image. Therefore, our final loss for DGS is as follows:
$\mathcal{L}_{\text{DGS}}=\mathcal{L}_{\text{AEGS}}+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}}=(1-\lambda)\mathcal{L}_{1}(\mathbf{A}_{i}(\mathcal{I}^{r}_{i}),\mathcal{I}_{i})+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}},$ (3)
where $\lambda_{{\text{reg}_{o}}}$ is the hyper-parameter for offset regularization.
@@ 5 Experiments
Figure 5: Qualitative comparison with NF-LDM (figure from Kim et al. (2023)) and original 3DGS (on the same generated video). Panoramas for GS are transformed and stitched from perspective cameras with $90^{\circ}$ FOV. Views in the last row are rendered with camera rigs different from the nuScenes dataset.
@@ 5.1 Experimental Setup
Dataset. We test our MagicDrive3D using the nuScenes dataset (Caesar et al., 2020), which is commonly used for generating and reconstructing street views (Gao et al., 2024; Wen et al., 2023; Wang et al., 2023b; Xie et al., 2023). The official configuration is followed, using 700 street-view video clips of approximately 20s each for training and another 150 clips for validation. For semantics in control signals, we follow Gao et al. (2024), using 10 object classes and 8 road classes.
@@
Metrics and Settings. MagicDrive3D is primarily evaluated using the Fréchet Inception Distance (FID) by rendering novel views unseen in the dataset and comparing their FID with real images. In addition, the method’s video generation ability is evaluated using Fréchet Video Distance (FVD), and its reconstruction performance is assessed using L1, PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018). For reconstruction evaluation, two testing scenarios are employed: 360, where all six views from $t=9$ are reserved for testing the reconstruction in the canonical space; and vary-t, where one view is randomly sampled from different $t$ to assess long-range reconstruction ability through $t$ in the canonical space (as shown in Table 1).
Implementation. For video generation, we train our generator based on the pre-trained street view image generation model from Gao et al. (2024). By adding the proposed relative pose control, we train 4 epochs (77040 steps) on the nuScenes training set with a learning rate of $8e^{-5}$. We follow the settings for 7-frame videos described in Gao et al. (2024), using $224\times 400$ for each view but extending to $T=16$ frames. Consequently, for reconstruction, we select $t=8$ as the canonical space. Except we change the first 500 steps to optimize $(s_{c,t},b_{c,t})$ for each view and $\lambda_{\text{reg}_{o}}=1.0$, other settings are the same as 3DGS. More details can be found in Appendix A.
5.2 Main Results
Table 2: Generation quality evaluation. All validation scenes from the nuScenes dataset are adopted. We use all generated views for reconstruction. Novel views adopt camera poses different from the nuScenes.
@@ 5.2 Main Results
Generation Quality. As shown in Table 2, the evaluation of generation quality involves two aspects. Firstly, the quality of video generation is assessed using extended 16-frame MagicDrive (Gao et al., 2024) as a baseline. Despite minor improvement in single-frame quality based on FID, MagicDrive3D substantially enhances video quality (as evidenced by FVD), demonstrating the efficacy of the proposed relative camera pose embedding in enhancing temporal consistency. Secondly, the image quality of renderings from the generated scene is evaluated using FID. We also include qualitative comparisons in Figure 5. Compared to 3DGS, our enhanced DGS significantly enhances visual quality in reconstructing contents, particularly in unseen novel views. More qualitative comparison can be found in Appendix C.
Reconstruction Quality. Our enhanced DGS, as a reconstruction method, is further evaluated by comparing renderings with ground truth (GT) images. Here, the generated views from the video generator are treated as GT. We employ two settings per Table 1, with results displayed in Table 3. As per all metrics, our enhanced DGS not only improves reconstruction quality for training views but also drastically enhances quality for testing views, compared to 3DGS. We include comparison with 4DGS (Wu et al., 2024) in Appendix B.
Table 3: Reconstruction quality evaluation. We random sample 100 scenes from the nuScenes validation set for evaluation. “cc” refers to color correction from Barron et al. (2022). Although 3DGS does not consider appearance differences, we apply “cc” to it for fair comparisons.
@@
5.3 Ablation Study
setting method L1 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
vary-t 3DGS 0.0799 19.1387 0.4814 0.4697
@@ 5.3 Ablation Study
Ablation on Offset Choice for Deformable GS. In addition to the overall module ablation, we observe that for Deformable GS, beyond the Gaussians’ center coordinates, their attributes (including anisotropic covariance, opacity, and harmonic coefficients) can also be utilized to address local inconsistencies. To verify the effects of different choices, we randomly select 10 scenes from the nuScenes validation set for experimentation. As shown in Table 5, the results obtained by adding offsets to the center coordinates (xyz) are the best. This aligns with our observation that local inconsistencies in the generated views occur primarily in the shape of objects, and thus, xyz displacements can most effectively resolve these inconsistencies.
Table 5: Ablating comparison with offsets on anisotropic covariance (Cov.), opacity, and harmonic coefficients (Features) properties in GS. We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
Methods L1 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS 0.0733 19.7514 0.5210 0.4496
Features offset 0.0624 20.9882 0.5940 0.3463
Cov. offset 0.0632 20.8133 0.5854 0.3626
Opacity offset 0.0656 20.5332 0.5733 0.3845
Ours (xyz offset) 0.0546 21.9428 0.6288 0.2759
Table 6: MagicDrive3D improves the viewpoint robustness (Klinghoffer et al., 2023) of CVT (Zhou & Krähenbühl, 2022). All results are mIoU for BEV segmentation. Colors highlight the differences with baseline. The best results are in bold.
@@ 5.4 Application
Training Support for Perception Tasks. We demonstrate an application wherein street scene generation serves as a data engine for perception tasks, leveraging the advantage of any-view rendering to improve viewpoint robustness (Klinghoffer et al., 2023). We employ CVT (Zhou & Krähenbühl, 2022) and the BEV segmentation task following the evaluation protocols of Zhou & Krähenbühl (2022); Gao et al. (2024). By incorporating 4 different rigs on the FRONT camera and adding rendered views for training, the negative impact from viewpoint changes is alleviated (Table 6), exemplifying the utility of street scene generation in training perception tasks.
Render Object-level Dynamic. MagicDrive3D generates 3DGS representations of scenes, thereby enabling applications for scene editing. Our approach employs metric scale modeling, ensuring that the editing of scenes corresponds accurately to real-world physical distances. As demonstrated in Figure 7, we segmented the generated GS and relocated the object on the right. The resulting scene GS supports rendering effectively.
(a) Original view.
(b) Move forward 0.1m.
(c) Move forward 1m.
(d) Move forward 1.6m.
@@ 6 Conclusion and Discussion
This paper introduces MagicDrive3D, a unique 3D street scene generation framework that integrates geometry-free view synthesis and geometry-focused 3D representations. MagicDrive3D significantly reduces data requirements, enabling training on typical autonomous driving datasets, such as nuScenes. Within the generation-reconstruction pipeline, MagicDrive3D employs a video generation model to enhance inter-frame consistency, while the enhanced deformable GS improves reconstruction quality from generated views. Comprehensive experiments demonstrate that MagicDrive3D can produce high-quality 3D street scenes with multi-level controls. Additionally, we show that scene generation can serve as a data engine for perception tasks such as BEV segmentation.
Limitation and Future Work. As a data-centric method, MagicDrive3D sometimes struggles to generate complex objects like pedestrians, whose appearances are intricate. Additionally, areas with much texture detail (e.g., road fences) or small spatial features (e.g., light poles) are occasionally poorly generated due to limitations in the reconstruction method. Future work may focus on addressing these challenges and further improving the quality and robustness of generated 3D scenes.
APPENDIX
Appendix A More Implementation Details
For the monocular depth model, we use ZoeDepth (Bhat et al., 2023). Although it is trained for metric depth estimation, due to domain differences, raw estimation is not usable, as shown in Section 5.3 and Figure 3. Methodologically, MagicDrive3D does not rely on a specific depth estimation model. Better estimations can further improve our scene generation quality.
@@ Appendix A More Implementation Details
Appendix B More Reconstruction Baseline
Focusing on dynamic scenes, 4DGS (Wu et al., 2024) introduces comprehensive improvements over 3DGS and achieves notable results. Therefore, we replace deformable GS with 4DGS. As shown in Table 7, by incorporating non-rigid dynamics, 4DGS already performs better than 3DGS. However, in our task, 4DGS underperforms compared to our deformable GS. Based on the results in Table 5, we hypothesize that our reconstruction algorithm only needs to address local dynamics caused by content inconsistency. Allocating excessive dynamics at the representational level may hinder model convergence, thus 4DGS does not yield better results in our scenarios.
Table 7: Comparison with 4DGS (Wu et al., 2024). We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
@@ Appendix B More Reconstruction Baseline
Appendix C Comparison with Simple Baselines
(a) Conditional image to LucidDreamer (Chung et al., 2023)
(b) Scene generated by LucidDreamer (Chung et al., 2023), with text “A driving scene in the city from the front camera of the vehicle. A bus on the right side. There is a bridge overhead. There is a railing in the center of the left road. Some vehicles ahead
(c) Scene generated by WonderJourney (Yu et al., 2024). We set the focal length to be the same as the conditional image. WonderJourney cannot control road semantics (many objects are physically implausible) and fails in loop closure for 360 scene generation. Conditions are the same as Figure 8(b).
@@ Appendix C Comparison with Simple Baselines
Figure 8: Comparison with two baselines (LucidDreamer (Chung et al., 2023) and WonderJourney (Yu et al., 2024)) and direct stitching real images.
As shown in Figure 8, we further compare MagicDrive3D with two baselines, i.e., LucidDreamer (Chung et al., 2023) and WonderJourney (Yu et al., 2024). The former method has been proposed recently and takes text description as the only condition. Thus, it is hard to generate photo-realistic street scenes. When providing multi-view video frames from nuScenes with known camera poses, their pipeline fails to reconstruct. We suppose the reason is limited overlaps and errors from depth estimation. As suggested by the released code, we changed the image generation model to lllyasviel/control_v11p_sd15_inpaint for inpainting by providing a nuScenes image, i.e., Figure 8(a). However, due to the lack of controllability, the results from LucidDream (e.g., Figure 8(b)) are unsatisfactory. On the other hand, due to the lack of control over objects within the scene, WonderJourney struggles to generate coherent scenes. Inpainting-based methods like the two above exhibit a pronounced sense of patches and face significant challenges in achieving 360 coverage.
@@
Appendix D Implementation Detail of Appearance Embedding
We show in Figure 9 the detailed architecture of the CNN used in our appearance modeling. The AE map is 32×32$\times$32 × smaller than the input image to reduce the computational cost. Hence, we first downsample the input image by 32×32$\times$32 ×. Then, we use $3\times 3$ convolution for feature extraction and pixel shuffle for upsampling. Each convolution layer is activated by ReLU.
Figure 9: The CNN architecture of appearance modeling, as introduced in Section 4.3.
Appendix E Broader Impacts

### v3

@@ 4.3 Enhanced GS for Generated Content
Consequently, $\boldsymbol{\mu}^{o}_{p}(t)$ can manage the local dynamics driven by pixel-level disparities, while $\boldsymbol{\mu}$ focuses on the global geometric correlations. It ensures the quality of scene reconstruction by leveraging consistent parts across different viewpoints, simultaneously eliminating artifacts. Besides, with the analytical gradient w.r.t. $SE(3)$ pose of cameras [17], we also make the camera pose optimizable in the final few steps of GS iterations, which helps to mitigate the local dynamic from camera poses.
@@
Figure 5: Qualitative comparison with NF-LDM [13] and original 3DGS (on the same generated video as ours). Panoramas for GS are transformed and stitched from perspective cameras with $90^{\circ}$ FOV. Views in the last row are rendered with unseen camera rigs of nuScenes.
Loss: Aligning Exposure with Appearance Modeling. Typical street view datasets are collected with multiple cameras, which capture views independently through auto-exposure and auto-white-balance [4]. Since the video generation is optimized to match the original data distribution, the differences from different cameras also exist in the generated data. The appearance differences are well-known issues for in-the-wild reconstruction [16]. In this paper, we propose a dedicated appearance modeling technique for GS representation.
We hypothesize that the disparity between different views can be represented by affine transformations $\mathbf{A}_{i}(\cdot)$ for $i$-th camera view. An Appearance Embedding (AE) map $\mathbf{e}_{i}\in\mathbb{R}^{w_{e}\times h_{e}\times c_{e}}$ is allocated for each view, and a Convolutional Neural Network (CNN) is utilized to approximate this transformation matrix $w_{\mathbf{A}}\in\mathbb{R}^{w\times h\times 3}$ (Appendix F contains more details). The final computation of the pixel-wise $\ell_{1}$ loss is conducted using the transformed image. Therefore, our final loss for DGS is as Equation 3, where $\lambda_{{\text{reg}_{o}}}$ is the hyper-parameter for offset regularization.
$\begin{split}\mathcal{L}_{\text{DGS}}&=\mathcal{L}_{\text{AEGS}}+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}}\\ &=(1-\lambda)\mathcal{L}_{1}(\mathbf{A}_{i}(\mathcal{I}^{r}_{i}),\mathcal{I}_{i})\\ &+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}}\end{split}$ (3)
@@ 5.1 Experimental Setup
Dataset. We test our MagicDrive3D using the nuScenes dataset [4], which is commonly used for generating and reconstructing street views [9, 32, 30, 34]. The official configuration is followed, using 700 street-view video clips of approximately 20s each for training and another 150 clips for validation. For semantics in control signals, we follow [9], using 10 object classes and 8 road classes.
@@
Metrics and Settings. MagicDrive3D is primarily evaluated using the Fréchet Inception Distance (FID) by rendering novel views unseen in the dataset and comparing their FID with real images. In addition, the method’s video generation ability is evaluated using Fréchet Video Distance (FVD), and its reconstruction performance is assessed using L1, PSNR, SSIM [31], and LPIPS [37]. For reconstruction evaluation, two testing scenarios are employed: 360, where all six views from $t=9$ are reserved for testing the reconstruction in the canonical space; and vary-t, where one view is randomly sampled from different $t$ to assess long-range reconstruction ability through $t$ in the canonical space (as shown in Table 1). More implementation details can be found in Appendix A.
@@
Table 2: Generation quality evaluation. We adopt all validation scenes from nuScenes. All generated views are used for reconstruction. Novel views are camera poses different from nuScenes.
@@
Table 3: Reconstruction quality evaluation. We randomly sample 100 scenes from the nuScenes val. set for evaluation. “cc” refers to color correction from [1]. Although 3DGS does not consider appearance differences, we apply “cc” to it for fair comparisons.
5.2 Main Results
Generation Quality. As shown in Table 2, the evaluation of generation quality involves two aspects. Firstly, the quality of video generation is assessed using extended 16-frame MagicDrive [9] as a baseline. Despite minor improvement in single-frame quality based on FID, MagicDrive3D substantially enhances video quality (as evidenced by FVD), demonstrating the efficacy of the proposed relative camera pose embedding in enhancing temporal consistency. Secondly, the image quality of renderings from the generated scene is evaluated using FID. We also include qualitative comparisons in Figure 5. Compared to 3DGS, our enhanced DGS significantly enhances visual quality in reconstructing contents, particularly in unseen novel views. More qualitative comparison can be found in Appendix E.
Reconstruction Quality. Our enhanced DGS, as a reconstruction method, is further evaluated by comparing renderings with ground truth (GT) images. Here, the generated views from the video generator are treated as GT. We employ two settings per Table 1, with results displayed in Table 3. As per all metrics, our enhanced DGS not only improves reconstruction quality for training views but also drastically enhances quality for testing views, compared to 3DGS. We include comparison and discussion with 4DGS [33] in Appendix C.
Methods L1 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
vary-t setting:
3DGS 0.0799 19.1387 0.4814 0.4697
@@ 5.2 Main Results
(a) Original view.
(b) Move forward 0.1m.
(c) Move forward 1m.
(d) Move forward 1.6m.
@@
Table 5: MagicDrive3D improves the viewpoint robustness [14] of CVT [39]. All results are mIoU for BEV segmentation. Colors highlight the differences with baseline. The best results are in bold.
5.3 Ablation Study
@@ 5.4 Application
Training Support for Perception Tasks. We demonstrate an application wherein street scene generation serves as a data engine for perception tasks, leveraging the advantage of any-view rendering ability of our method to improve viewpoint robustness for perception models [14]. We employ CVT [39] and the BEV segmentation task following the evaluation protocols of [39, 9]. By incorporating 4 different rigs on the FRONT camera and adding rendered views for training, the negative impact from viewpoint changes is alleviated (Table 5), exemplifying the utility of street scene generation in training perception tasks.
Render Object-level Dynamic. MagicDrive3D is capable of generating 3DGS representations of scenes, which facilitates various applications in scene editing. Our method utilizes metric scale modeling, guaranteeing that any modifications made to the scenes accurately reflect real-world physical distances. As illustrated in Figure 6, we segmented the generated GS and repositioned the object on the right. The resulting scene GS supports rendering effectively.
@@ 6 Conclusion and Discussion
This paper introduces MagicDrive3D, a unique 3D street scene generation framework that integrates geometry-free view synthesis and geometry-focused 3D representations. MagicDrive3D significantly reduces data requirements, enabling training on typical autonomous driving datasets, such as nuScenes. MagicDrive3D employs a video generation model to enhance inter-frame consistency, while the enhanced deformable GS improves reconstruction quality from generated views. Comprehensive experiments demonstrate that MagicDrive3D can produce high-quality 3D street scenes with multi-level controls. Additionally, we show that scene generation can serve as a data engine for perception tasks such as BEV segmentation.
Supplementary Material (Appendix)
Please find the video visualizations on our project website: https://flymin.github.io/magicdrive3d/
A More Implementation Details
For video generation, we train our generator based on the pre-trained street view image generation model from [9]. By adding the proposed relative pose control, we train 4 epochs (77040 steps) on the nuScenes training set with a learning rate of $8e^{-5}$. We follow the settings for 7-frame videos described in [9], using $224\times 400$ for each view but extending to $T=16$ frames. Consequently, for reconstruction, we select $t=8$ as the canonical space. Except we change the first 500 steps to optimize $(s_{c,t},b_{c,t})$ for each view and $\lambda_{\text{reg}_{o}}=1.0$, other settings are the same as 3DGS.
For the monocular depth model, we use ZoeDepth [3]. Although it is trained for metric depth estimation, due to domain differences, raw estimation is not usable, as shown in Section 5.3 and Figure 3. Methodologically, MagicDrive3D does not rely on a specific depth estimation model. Better estimations can further improve our scene generation quality.
@@ A More Implementation Details
B More Ablation Study
Ablation on Offset Choice for Deformable GS. In addition to the overall module ablation, we observe that for Deformable GS, beyond the Gaussians’ center coordinates, their attributes (including anisotropic covariance, opacity, and harmonic coefficients) can also be utilized to address local inconsistencies. To verify the effects of different choices, we randomly select 10 scenes from the nuScenes validation set for experimentation. As shown in Table I, the results obtained by adding offsets to the center coordinates (xyz) are the best. This aligns with our observation that local inconsistencies in the generated views occur primarily in the shape of objects, and thus, xyz displacements can most effectively resolve these inconsistencies.
Methods L1 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS 0.0733 19.7514 0.5210 0.4496
Features offset 0.0624 20.9882 0.5940 0.3463
Cov. offset 0.0632 20.8133 0.5854 0.3626
Opacity offset 0.0656 20.5332 0.5733 0.3845
Ours (xyz offset) 0.0546 21.9428 0.6288 0.2759
Table I: Ablating comparison with offsets on anisotropic covariance (Cov.), opacity, and harmonic coefficients (Features) properties in GS. We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
C More Reconstruction Baseline
Focusing on dynamic scenes, 4DGS [33] introduces comprehensive improvements over 3DGS and achieves notable results. Therefore, we replace deformable GS with 4DGS. As shown in Table II, by incorporating non-rigid dynamics, 4DGS already performs better than 3DGS. However, in our task, 4DGS underperforms compared to our deformable GS. Based on the results in Table I, we hypothesize that our reconstruction algorithm only needs to address local dynamics caused by content inconsistency. Allocating excessive dynamics at the representational level may hinder model convergence, thus 4DGS does not yield better results in our scenarios.
@@ C More Reconstruction Baseline
Table II: Comparison with 4DGS [33]. We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
D Limitation and Future Work
As a data-centric method, MagicDrive3D sometimes struggles to generate complex objects like pedestrians, whose appearances are intricate. Additionally, areas with much texture detail (e.g., road fences) or small spatial features (e.g., light poles) are occasionally poorly generated due to limitations in the reconstruction method. Future work may focus on addressing these challenges and further improving the quality and robustness of generated 3D scenes.
E Comparison with Simple Baselines
(a) Conditional image to LucidDreamer [6]
(b) Scene generated by LucidDreamer [6], with text “A driving scene in the city from the front camera of the vehicle. A bus on the right side. There is a bridge overhead. There is a railing in the center of the left road. Some vehicles ahead
(c) Scene generated by WonderJourney [36]. We set the focal length to be the same as the conditional image. WonderJourney cannot control road semantics (many objects are physically implausible) and fails in loop closure for 360 scene generation. Conditions are the same as Figure 1(b).
@@ E Comparison with Simple Baselines
Figure I: Comparison with two baselines (LucidDreamer [6] and WonderJourney [36]) and direct stitching real images.
As shown in Figure I, we further compare MagicDrive3D with two baselines, i.e., LucidDreamer [6] and WonderJourney [36]. The former method has been proposed recently and takes text description as the only condition. Thus, it is hard to generate photo-realistic street scenes. When providing multi-view video frames from nuScenes with known camera poses, their pipeline fails to reconstruct. We suppose the reason is limited overlaps and errors from depth estimation. As suggested by the released code, we changed the image generation model to lllyasviel/control_v11p_sd15_inpaint for inpainting by providing a nuScenes image, i.e., Figure 1(a). However, due to the lack of controllability, the results from LucidDream (e.g., Figure 1(b)) are unsatisfactory. On the other hand, due to the lack of control over objects within the scene, WonderJourney struggles to generate coherent scenes. Inpainting-based methods like the two above exhibit a pronounced sense of patches and face significant challenges in achieving 360 coverage.
@@
F Implementation Detail of Appearance Embedding
We show in Figure II the detailed architecture of the CNN used in our appearance modeling. The AE map is 32×32$\times$32 × smaller than the input image to reduce the computational cost. Hence, we first downsample the input image by 32×32$\times$32 ×. Then, we use $3\times 3$ convolution for feature extraction and pixel shuffle for upsampling. Each convolution layer is activated by ReLU.
Figure II: The CNN architecture of appearance modeling, as introduced in Section 4.3.
G Broader Impacts